# Machine-Learning Characterization of Intermittency in Plasma Turbulence: Single and Double Sheet Structures

### v2

@@ 5 Discussion
The formation of the single sheets is governed by the dynamics of the coherent structures in the turbulent flow (e.g., Zhou et al., 2023). First, we note that MHD turbulence tends to generate and sustain magnetic flux ropes (helical magnetic field structures with an electric current flowing along $\mathbf{B}$). The dynamics between such structures follow from basic electrodynamics: flux ropes with the same polarity attract; opposite polarities repel each other. Mergers of the same-polarity flux ropes form a reversal in the $\delta\mathbf{B}\perp\mathbf{B}_{0}$ magnetic field components; such region naturally induces a spatio-temporal current sheet. Such single sheets are prone to magnetic reconnection (e.g., Priest & Forbes, 2000; Zweibel & Yamada, 2009) and their lifetime is regulated by the tearing-instability (e.g., Comisso & Sironi, 2019) (or their environment). They are well-known sites of plasma energization (Sironi & Spitkovsky, 2014; Nättilä & Beloborodov, 2021).
@@
Importantly, our analysis also demonstrates that plasma turbulence tends to host a second type of current sheets, which has been largely overlooked so far. These double sheets have been seen in previous numerical studies (e.g., Servidio et al., 2009; Zhou et al., 2023) and observations of the Earth’s magnetosphere (e.g., Ergun et al., 2009); however, their statistics has not been analyzed nor has their importance been highlighted in the past. The physical origin of the double current sheets remains uncertain. Most likely, the double sheets arise from the nonlinear interactions between Alfvén wave packets (Howes & Nielson, 2013; Howes, 2016; Verniero et al., 2018). More specifically, we observe that the double current sheets are formed where magnetic structures of opposite polarities (i.e., rotational directions) are interacting. However, even if their origin is not yet understood, their existence is important for future analysis of turbulence. For example, it is not yet understood what scatters the energetic cosmic rays in the inter- and intra-galactic plasmas: one hypothesis is the intermittent structures in the turbulence (Kempski et al., 2023; Lemoine, 2023). A drastic difference in the shape of the (magnetic) structures can then lead to a significant difference in the scattering properties. Lastly, the two intermittent structures also have differing dissipation profiles, which need to be accounted for in turbulence sub-grid models.
@@
The authors thank Shirley Ho and Kaze Wong for useful discussions. TH acknowledges support from a pre-doctoral program at the Flatiron Institute. Research at the Flatiron Institute is supported by the Simons Foundation. JN is supported by an ERC grant (ILLUMINATOR, 101114623). JD is supported by NASA through the NASA Hubble Fellowship grant HST-HF2-51552.001-A, awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. LS acknowledges support from DoE Early Career Award DE-SC0023015. This work was supported by a grant from the Simons Foundation (MP-SCMPS-00001470) to LS, and facilitated by Multimessenger Plasma Physics Center (MPPC), NSF grant PHY-2206609 to LS. Simons Foundation is acknowledged for computational support.
@@ Appendix A SOM segmentation method
We develop a custom-made SOM implementation, aweSOM (Ha et al., 2024, submitted). This Python package employs the same technique as similar Python-based SOM implementations. Specifically, we base aweSOM on POPSOM [^2] https://github.com/njali2001/popsom. We introduce optimizations to improve analysis speed, enabling it to handle complex, high-dimensional physical data. Below, we briefly describe aweSOM’s approach to learning the magnetized plasma turbulence.
@@
Lastly, the cells in the simulation are mapped to the nearest node in the lattice, each of which has been assigned a cluster label. This label is then transferred to the corresponding input vector, resulting in visualization of the clustering in the input space. An example of this mapping is shown in the bottom panel of Figure 1.
@@ Appendix E Statistical Distributions of SOM clusters
Figure 6: Feature-to-feature scatter plots of SOM clusters from Figure 2. For clarity, cluster 6 has been excluded, and a random sample of 50000 points ($\sim 0.025\,L^{3}$) is drawn for each plot. Each data point is color-coded by its assigned cluster label.
@@
Figure 6 shows scatter plots between pairs of features used. A random sample of $5\times 10^{4}$ data points is used, to improve the clarity of the plots. The data points are quite clearly separated in a multi-dimensional space. For instance, clusters 0 (blue) and 4 (purple), as well as clusters 1 (orange) and 2 (green), are separated along the $j_{\parallel}$$=0$ line. As we established in the main text, these clusters are either parts of the double sheet structures (clusters 1 and 2), or they are individual current sheets (clusters 0 and 4). Another correlation that we can see is in the $e_{\parallel}$ plots, where clusters 3 (red) and 5 (brown) are on opposite tail ends of the distribution. We also observe that the double current sheets tend to have markedly stronger $b_{\perp}$ than the rest of the data, pointing to the prominent peak in $B_{p1}$ seen in Figure 5.
@@ F.1 Reconnecting Harris current sheet
Figure 8: Left: $xy$-slice of current density along z-axis ($J_{z}$, top) vs. aweSOM clustering result (bottom) at $z=-40c/\omega_{\rm p}$ for the Harris current sheet simulation. Streamlines in the top panel show in-plane magnetic field lines. Contours in both panels show regions where $J_{z}>2J_{\rm rms}$, representing the boundary of the Harris current sheet. Right: $xy$-slice of $j_{\parallel}$ (top) vs. aweSOM clustering result (bottom) at $z=120c/\omega_{\rm p}$. Streamlines in the top panel show the in-plane magnetic field, $B_{xy}$. Red and blue contours in the bottom panel show locations where $j_{\parallel}<-3\,j_{\parallel,\rm rms}$, and where $j_{\parallel}<-3\,j_{\parallel,\rm rms}$, respectively. The color of the clusters have the same physical meaning as in Figure 3
The left side of Figure 8 shows a $xy$-slice of an aweSOM clustering result of the Harris sheet. In this idealized simulation, cluster 2 (green) has an almost-one-to-one correspondence with a current density threshold of $J_{z}>2J_{\rm rms}$, commonly used as a proxy for identifying current sheets in simulations (Zhdankin et al., 2013). Cluster 0 (blue) accounts for the rest of the background plasma, and cluster 1 (orange) are locations with strong $e_{\parallel}$.
@@ F.2 Driven turbulence
We also use snapshots from a continuously driven turbulence simulation (Nättilä, 2024). Similar to the freely-evolving turbulence case, the domain is a triply periodic cubic box of length $L=640c/\omega_{\rm p}$, which contains a neutral $e^{\pm}$-pair plasma with 8 particles per cell per species. The domain has a magnetization of $\sigma=10$ and is initialized with a uniform magnetic field $\mathbf{B}_{0}=B_{0}\hat{\mathbf{z}}$. The turbulence is continuously excited by driving an external current with an oscillating Langevin antenna formalism (TenBarge et al., 2014). The antenna has the following properties: driving scale $l_{0}=L/2$, amplitude $\delta B/B_{0}=0.8$, frequency $\omega_{\rm ant}/\omega_{0}=0.8$ (where $\omega_{0}=2\pi/t_{0}$ is the eddy-turnover frequency, and $t_{0}=l_{0}/v_{\rm A}$ is the eddy-turnover time), and a decorrelation time of $\omega_{\rm dec}/\omega_{0}=0.6$.
@@
Notably, the volume-filling fraction of the sheets is much smaller than that found in the freely-evolving turbulence simulation, $n_{\rm fill,0+1+4}\approx 1.37\%$. Given the similar (down-sampled) resolution of both the decaying turbulence and driven turbulence box, such a discrepancy in volume-filling fractions may be explained by the different modes of perturbation, or by the size of the turbulence driver ($l_{0}=L/2$, vs. $L/4$ in the decaying turbulence simulation). Alternatively, the difference seen here could be due to a transient turbulent effect such as a flare, that is inadequately captured by individual snapshots.

### v3

@@ 5 Discussion
The formation of the single sheets is governed by the dynamics of the coherent structures in the turbulent flow (e.g., Zhou et al., 2023). First, we note that MHD turbulence tends to generate and sustain magnetic flux ropes (helical magnetic field structures with an electric current flowing along $\mathbf{B}$). The dynamics between such structures follow from basic electrodynamics: flux ropes with the same polarity attract; opposite polarities repel each other. Mergers of the same-polarity flux ropes form a reversal in the $\delta\mathbf{B}\perp\mathbf{B}_{0}$ magnetic field components; such region naturally induces a spatio-temporal current sheet. Such single sheets are prone to magnetic reconnection (e.g., Priest & Forbes, 2000; Zweibel & Yamada, 2009) and their lifetime is regulated by the tearing-instability (e.g., Comisso & Sironi, 2019) (or their environment). They are well-known sites of plasma energization (Sironi & Spitkovsky, 2014; Comisso & Sironi, 2018).
@@
Importantly, our analysis also demonstrates that plasma turbulence tends to host a second type of current sheets, which has been largely overlooked so far. These double sheets have been seen in previous numerical studies (e.g., Servidio et al., 2009; Zhou et al., 2023) and observations of the Earth’s magnetosphere (e.g., Ergun et al., 2009); however, their statistics has not been analyzed nor has their importance been highlighted in the past. The physical origin of the double current sheets remains uncertain. Most likely, the double sheets arise from the nonlinear interactions between Alfvén wave packets (Howes & Nielson, 2013; Howes, 2016; Verniero et al., 2018). In a freely-evolving relativistic pair plasma turbulence, we observe that the double current sheets are formed where magnetic structures of opposite polarities (i.e., rotational directions) are interacting. However, even if their origin is not yet understood, their existence is important for future analysis of turbulence. For example, it is not yet understood what scatters the energetic cosmic rays in the inter- and intra-galactic plasmas: one hypothesis is the intermittent structures in the turbulence (Kempski et al., 2023; Lemoine, 2023). A drastic difference in the shape of the (magnetic) structures can then lead to a significant difference in the scattering properties. Lastly, the two intermittent structures also have differing dissipation profiles, which need to be explicitly accounted for in future turbulence sub-grid models that make assumptions about the local heating and particle acceleration. For example, Howes et al. (2018) and Nättilä & Beloborodov (2022) found that Landau damping plays an important role in the energy dissipation of kinetic-Alfvén-wave turbulence. Future studies on the stability of current sheets should also investigate the dissipation profile of these sheets, and whether they are analogous to the “double tearing mode” current sheets in the literature (e.g., Pritchett et al., 1980; Baty et al., 2013).
@@
We would like to thank the anonymous reviewer for their constructive feedback which improves the clarity of this paper. The authors thank Shirley Ho and Kaze Wong for useful discussions. TH acknowledges support from a pre-doctoral program at the Flatiron Institute. Research at the Flatiron Institute is supported by the Simons Foundation. JN is supported by an ERC grant (ILLUMINATOR, 101114623). JD is supported by NASA through the NASA Hubble Fellowship grant HST-HF2-51552.001-A, awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. LS acknowledges support from DoE Early Career Award DE-SC0023015. This work was supported by a grant from the Simons Foundation (MP-SCMPS-00001470) to LS, and facilitated by Multimessenger Plasma Physics Center (MPPC), NSF grant PHY-2206609 to LS. LS is also supported by NSF grant PHY-2409223. Simons Foundation is acknowledged for computational support.
@@ Appendix A SOM segmentation method
We develop a custom-made SOM implementation, aweSOM (Ha et al., 2025). This Python package employs the same technique as similar Python-based SOM implementations. Specifically, we base aweSOM on POPSOM [^2] https://github.com/njali2001/popsom. We introduce optimizations to improve analysis speed, enabling it to handle complex, high-dimensional physical data. Below, we briefly describe aweSOM’s approach to learning the magnetized plasma turbulence.
@@
Lastly, the cells in the simulation are mapped to the nearest node in the lattice, each of which has been assigned a cluster label. This label is then transferred to the corresponding input vector, resulting in visualization of the clustering in the input space. An example of this mapping is shown in panel c) of Figure 1.
@@ Appendix E Statistical Distributions of SOM clusters
Figure 6: Feature-to-feature scatter plots of SOM clusters from Figure 2. For clarity, cluster 6 has been excluded, and a random sample of $10^{5}$ points ($\sim 5\%\,L^{3}$) is drawn for each plot. Each data point is color-coded by its assigned cluster label.
Figure 7: Scatter plots of SOM clusters (in the background; similar to Figure 6) and projections of the SOM lattice (in the foreground), where each square is a node, and the stars are locations of the centroids of each cluster. The gray lines represent connected nodes along the $X$-axis in Figure 2. The two black curves show the horizontal boundaries of the map with $Y=0$ and $Y=36$.
@@
Figure 6 shows scatter plots between pairs of features used. A random sample of $10^{5}$ data points is used, and cluster 6 (background) is omitted, to improve the clarity of the plots. The data points are quite clearly separated in a multi-dimensional space. For instance, clusters 0 (blue) and 4 (purple), as well as clusters 1 (orange) and 2 (green), are separated along the $j_{\parallel}$$=0$ line. As we established in the main text, these clusters are either parts of the double sheet structures (clusters 1 and 2), or they are individual current sheets (clusters 0 and 4). Another correlation that we can see is in the $e_{\parallel}$ plots, where clusters 3 (red) and 5 (brown) are on opposite tail ends of the distribution. We also observe that the double current sheets tend to have markedly stronger $b_{\perp}$ than the rest of the data, pointing to the prominent peak in $B_{p1}$ seen in Figure 5. Figure 7 shows the same scatter plots as Figure 6, but with the projections of the SOM lattice plotted in the foreground, showing the non-linear topology of the lattice. It can be used to read the non-linear combination of the feature-space parameters that the map uses for the clustering: the major axis (i.e., slices along $X$) is mainly oriented along the $j_{a}$ feature but also captures variations in $e_{\parallel}$ in the top part of the map (for $Y\gtrsim 30$). The minor axis (i.e., slices along $Y$) is somewhat oriented along $j_{\parallel}$ and $b_{\perp}$. However, we emphasize that the real map is composed of a non-linear combination of the features and is, therefore, highly tangled in the feature space.
@@ F.1 Reconnecting Harris current sheet
Figure 9: $xy$-slice of current density along z-axis ($J_{z}$, top) vs. aweSOM clustering result (bottom) at $z=-40c/\omega_{\rm p}$ for the Harris current sheet simulation. Streamlines in the top panel show in-plane magnetic field lines. Contours in both panels show regions where $J_{z}>2J_{\rm rms}$, representing the boundary of the Harris current sheet.
@@
Figure 9 shows a $xy$-slice of an aweSOM clustering result of the Harris sheet. In this idealized simulation, cluster 2 (green) has an almost-one-to-one correspondence with a current density threshold of $J_{z}>2J_{\rm rms}$, commonly used as a proxy for identifying current sheets in simulations (Zhdankin et al., 2013). Cluster 0 (blue) accounts for the rest of the background plasma, and cluster 1 (orange) are locations with strong $e_{\parallel}$.
@@ F.2 Driven turbulence
Figure 10: $xy$-slice of $j_{\parallel}$ (top) vs. aweSOM clustering result (bottom) at $z=120c/\omega_{\rm p}$. Streamlines in the top panel show the in-plane magnetic field, $B_{xy}$. Red and blue contours in the bottom panel show locations where $j_{\parallel}<-3\,j_{\parallel,\rm rms}$, and where $j_{\parallel}<-3\,j_{\parallel,\rm rms}$, respectively. The color of the clusters have the same physical meaning as in Figure 3.
We also use snapshots from a continuously driven turbulence simulation (Nättilä, 2024). Similar to the freely-evolving turbulence case, the domain is a triply periodic cubic box of length $L=640c/\omega_{\rm p}$, which contains a neutral $e^{\pm}$-pair plasma with 8 particles per cell per species. The domain has a magnetization of $\sigma=10$ and is initialized with a uniform magnetic field $\mathbf{B}_{0}=B_{0}\hat{\mathbf{z}}$. The turbulence is continuously excited by driving an external current with an oscillating Langevin antenna formalism (TenBarge et al., 2014). The antenna has the following properties: driving scale $l_{0}=L/2$, amplitude $\delta B/B_{0}=0.8$, frequency $\omega_{\rm ant}/\omega_{0}=0.8$ (where $\omega_{0}=2\pi/t_{0}$ is the eddy-turnover frequency, and $t_{0}=l_{0}/v_{\rm A}$ is the eddy-turnover time), and a decorrelation time of $\omega_{\rm dec}/\omega_{0}=0.6$. The SOM presented below was trained on one snapshot of the simulation, taken at $t\approx 1.5\;l_{0}/c$, which exhibits a magnetic power spectrum $\propto k_{\perp}^{\alpha}$ with a slope of $\alpha\approx-2$.
@@
Notably, the volume-filling fraction of the sheets is much smaller than that found in the freely-evolving turbulence simulation, $n_{\rm fill,0+1+4}\approx 1.37\%$. Given the similar (down-sampled) resolution of both the decaying turbulence and driven turbulence box, such a discrepancy in volume-filling fractions may be explained by the different modes of perturbation, or by the size of the turbulence driver ($l_{0}=L/2$, vs. $L/4$ in the decaying turbulence simulation). Alternatively, the difference seen here could be due to a transient turbulent effect such as a flare, that is inadequately captured by individual snapshots. The largest current sheets in the continuously driven turbulent domain could be sourced by the external driving mechanism, rather than the non-linear turbulent cascade, and, hence, we do not expect the double sheets to be as prominent in this simulation compared to the fiducial case presented in the main text. Nevertheless, the analysis shows qualitatively similar results to the freely-evolving turbulence simulations.